OpenStudy (anonymous):

solve y = -0.2x^2 + 4.8x -10.8 using the quadratic formula

OpenStudy (saifoo.khan):

\[x = \frac{-(4.8) \pm \sqrt{4.8^2-4(-0.2)(-10.8)} }{2(-0.2)}\]

OpenStudy (anonymous):

So would the answer come out to be 12√14.4?

OpenStudy (saifoo.khan):

you are close.

OpenStudy (saifoo.khan):

\[3(4- \sqrt{10}), 3(4+\sqrt{10})\]

OpenStudy (anonymous):

how did you get that?

OpenStudy (saifoo.khan):

calc.

OpenStudy (saifoo.khan):

you have to solve the formula to get it.

OpenStudy (anonymous):

ok let me see if I can get it. Did you use a GCF?

OpenStudy (anonymous):

Or did you do something to the decimals?

OpenStudy (saifoo.khan):

we can convert that to decimals as well

OpenStudy (anonymous):

Because I redid solving the equation but I get the same thing :/

OpenStudy (saifoo.khan):

can u show me your working?

OpenStudy (anonymous):

\[-4.8 \pm \sqrt{23.04 - 8.64} \]

OpenStudy (anonymous):

/-0.4 (This is the denominator )